Find the area between y = x^3−81 x and the x-axis
Answer by Alan3354(69443)

Find the area between y = x^3 - 81 x and the x-axis
---------------
x-ints at -9, 0 & +9
--------
f(x) = x^3 - 81x
INT(x) = x^4/4 - 81x^2/2 (Ignore the C)
-----
INT(9) = 6561/4 - 6561/2 = -6561/4
INT(0) = 0
Area above the x-axis = 6561/4 sq units
Area below = 6561/4
Total bounded area = 6561/2 sq units